My teacher taught me to use a table but I did and I still don't understand. If anyone could help me I would really appreciate it!
A swimmer can swim 4km with the current in 24 minutes. The same distance would take 40 minutes against the current. Find the rate of the current and the speed of the swimmer.
So first I made the table using the formula d = r (t) and so my two equations are:
4 = (r-c)(40)
4 = (r+c)(24)
And then I added them together getting: 8 = 64r - 16c. I know I got this wrong but I don't understand why or how. Could you walk me through all of the steps? Please & thank you! :)

Answer by mananth(16949)

A swimmer can swim 4km with the current in 24 minutes. The same distance would take 40 minutes against the current. Find the rate of the current and the speed of the swimmer.
So first I made the table using the formula d = r (t) and so my two equations are:
4 = (r-c)(40) This is where you haVE MADE A MISTAKE. 4 IS KM & 40 MINS. CONVERT TO HOUR. ELSE YOU WILL GET KM/MINUTE AS SPEED.
4 = (r+c)(24)
4 = (r-c)(2/3)
multiply by 3
3*4= 2(r-c)
/2
6=r-c...........................(1)
4 = (r+c)(2/5)
similarly
10=r+c.............................(2)
add (1)&(2)
you get 2r=16
/2
r=8 km/h
r+c= 10
so c= 2km/h
...
m.ananth@hotmail.ca